# Centre for Cyber-Physical Systems IISc, Bengaluru, India

### v1

@@ 2.4 Problem Formulation
The problem 2.13 can be reformulated as finding the $\delta$-ISS-CLF function that satisfies the conditions of Definition 2.7 under the existence of some forward invariant controller $g$.
@@ 3 Neural $\delta$-ISS Control Lyapunov Function
$\forall x,\hat{x}\in{\mathbf{X}}:$
$\quad-V(x,\hat{x})+\alpha_{1}(|x-\hat{x}|)\leq\eta,$ (6a)
$\quad V(x,\hat{x})-\alpha_{2}(|x-\hat{x}|)\leq\eta,$ (6b)
$\forall x,\hat{x}\in{\mathbf{X}},\forall w,\hat{w}\in{\mathbf{W}}:$
$\quad V(f(x,g(x,w)),f(\hat{x},g(\hat{x},\hat{w})))-V(x,\hat{x})+\alpha_{3}(|x-\hat{x}|)-\sigma(|w-\hat{w})|)\leq\eta,$ (6c)
$\forall x\in{\mathbf{X}},\forall w\in{\mathbf{W}}:$
$\quad h(f(x,g(x,w)))-\kappa(h(x))\leq\eta,$ (6d)
@@
Note that $V(\cdot,\cdot)\in\{V|V:{\mathbf{X}}\times{\mathbf{X}}\rightarrow{\mathbb{R}}_{0}^{+}\},g(\cdot,\cdot)\in\{g|g:{\mathbf{X}}\times{\mathbf{W}}\rightarrow{\mathbf{U}}\},\alpha_{1},\alpha_{2},\alpha_{3}\in\mathcal{K}_{\infty},\sigma\in\mathcal{K}$. If the optimal solution of ROP $\eta^{*}\leq 0$, under the constraints defined in (6) leads to the satisfaction of conditions of Definition 2.7 and 2.10, and thus the function $V(x,\hat{x})$ will be a valid $\delta$-ISS-CLF for the unknown system enforcing the system to be $\delta$-ISS under the forward invariant controller $g$.
@@
    The function $f$ is unknown, therefore we cannot directly incorporate conditions (6c) and (6d). So, solving the ROP becomes non-trivial.
@@
    The structures of the $\mathcal{K}_{\infty}$ functions $\alpha_{1},\alpha_{2},\alpha_{3}$ and class $\mathcal{K}$ function $\sigma$ are unknown. Also, the function $\kappa$ is unknown in condition (6d). Hence, the optimization problem is non-convex.
@@
To overcome the aforementioned challenges, the following subsections made some assumptions. To address challenge (C1), we parametrize $\delta$-ISS-CLF and the controller as feed-forward neural networks denoted by $V_{\theta,b}$ and $g_{\bar{\theta},\bar{b}}$, respectively, where $\theta,\bar{\theta}$ are weight matrices and $b,\bar{b}$ are bias vectors. The detailed structures of these neural networks are discussed in the next subsection.
3.1 Neural Network Structure
The $\delta$-ISS-CLF neural network consists of an input layer with $2n$ (two times the dimension of the system $n$) and an output layer with one neuron, reflecting the scalar nature of $\delta$-ISS-CLF. The network includes $l_{v}$ hidden layers with each hidden layer containing $h_{v}^{i},i\in[1;l_{v}]$ neurons, where both the values are arbitrarily chosen.
The activation function of all the layers except the output layer is chosen to be any slope-restricted nonlinear function $\varphi(\cdot)$ (for example, ReLU, Sigmoid, Tanh etc.). Hence the resulting neural network function is obtained by recursively applying the activation function as follows:
[…]
@@ 3.1 Neural Network Structure
3.2 Training Data Generation and Formal Correctness
To overcome the challenge (C2) in ROP, we have the following assumption:
@@ 3.2 Training Data Generation and Formal Correctness
We consider having access to the black box or simulator model of the system. Hence, given a state-input pair $(x,u)$ we will be able to generate the next state $f(x,u)$.
To alleviate the challenge (C3), we make another assumption, which is mentioned below.
@@
We assume that the class $\mathcal{K}_{\infty}$ functions $\alpha_{i},i\in\{1,2,3\}$ are of degree $\gamma_{i}$ with respect to $|x-\hat{x}|$ and the class $\mathcal{K}$ function $\sigma$ is of degree $\gamma_{w}$ with respect to $|w-\hat{w}|$, i.e., $\alpha_{i}(|x-\hat{x}|)=k_{i}|x-\hat{x}|^{\gamma_{i}}$, $i\in\{1,2,3\}$ and $\sigma(|w-\hat{w}|)=k_{w}|w-\hat{w}|^{\gamma_{w}}$, where $k:=[k_{1},k_{2},k_{3},k_{w}]$ and $\gamma:=[\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{w}]$ are user-defined parameters. We also consider $\kappa(h(x))=k_{h}h(x)$ where $k_{h}\in[0,1]$ is also a user-defined parameter.
Now, for training the neural $\delta$-ISS-CLF and the controller, one requires data from state space and input space. Here, we adopt a sampling-based approach to obtain samples from the compact state space ${\mathbf{X}}$ and the input space ${\mathbf{W}}$. We collect samples $x_{s}$ and $w_{p}$ from ${\mathbf{X}}$ and ${\mathbf{W}}$, where $s=[1;N],p=[1;M]$. Consider ball $B_{\varepsilon_{x}}(x_{s})$ and $B_{\varepsilon_{u}}(w_{p})$ around each sample $x_{s}$ and $w_{p}$, with radius $\varepsilon_{x}$ and $\varepsilon_{u}$, such that, ${\mathbf{X}}\subseteq\bigcup_{s=1}^{N}B_{\varepsilon_{x}}(x_{s})$ and ${\mathbf{W}}\subseteq\bigcup_{p=1}^{M}B_{\varepsilon_{u}}(w_{p})$ with :
$|x-x_{s}|$ $\leq\varepsilon_{x},\forall x\in B_{\varepsilon_{x}}(x_{s}),$ (7a)
$|w-w_{p}|$ $\leq\varepsilon_{u},\forall w\in B_{\varepsilon_{u}}(w_{p}).$ (7b)
@@
$\mathcal{X}$ =$=\{x_{s}:|x-x_{s}|\leq\varepsilon_{x},\forall x\in B_{\varepsilon_{x}}(x_{s}),\forall s\in[1;N]\},$ (8a)
${\mathcal{W}}$ =$=\{w_{p}:|w-w_{p}|\leq\varepsilon_{u},\forall w\in B_{\varepsilon_{u}}(w_{p}),\forall p\in[1;M]\}.$ (8b)
@@
$\forall x_{q},x_{r}\in\mathcal{X}:$
$\quad-V_{\theta,b}(x_{q},x_{r})+k_{1}|x_{q}-x_{r}|^{\gamma_{1}}\leq\eta,$ (9a)
$\quad V_{\theta,b}(x_{q},x_{r})-k_{2}|x_{q}-x_{r}|^{\gamma_{2}}\leq\eta,$ (9b)
$\forall x_{q},x_{r}\in\mathcal{X},\forall w_{q},w_{r}\in{\mathcal{W}}:$
$\quad V_{\theta,b}(f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q})),f(x_{r},g_{\bar{\theta},\bar{b}}(x_{r},w_{r})))-V_{\theta,b}(x_{q},x_{r})+k_{3}|x_{q}-x_{r}|^{\gamma_{3}}-k_{w}|w_{q}-w_{r}|^{\gamma_{w}}\leq\eta,$ (9c)
$\forall x_{q}\in\mathcal{X},w_{q}\in{\mathcal{W}}:$
$\quad h(f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q})))-k_{h}h(x_{q})\leq\eta.$ (9d)
Given the finite number of data samples, the SCP involves a finite set of constraints, making its solution computationally tractable. Let $\eta_{S}^{*}$ denote the optimal solution of the SCP. To demonstrate that the solution of the SCP is also a feasible solution for the proposed ROP, we impose the following assumptions regarding Lipschitz continuity:
@@
Since the model is unknown, these constants can be estimated following a similar procedure in [26, Algorithm 2].
@@
We assume that the candidate neural $\delta$-ISS-CLF is Lipschitz continuous with Lipschitz bound $\mathcal{L}_{L}$ with respect to $(x,\hat{x})$ over the set ${\mathbf{X}}$. Similarly, the controller neural network has a Lipschitz bound $\mathcal{L}_{C}$.
@@
The class $\mathcal{K}_{\infty}$ functions and the class $\mathcal{K}$ function of Definition 2.7 are Lipschitz continuous with Lipschitz constants $\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}$, and $\mathcal{L}_{w}$, respectively. One can estimate the values using the values of $k$ and $\gamma$. In addition, the Lipschitz constant $\mathcal{L}_{h}$ of the function $h$ is already predefined due to the known structure of $h$.
Under assumptions 3 and 4, the following theorem outlines the connection of the solution of SCP (9) to that of ROP (6), providing a formal guarantee to the obtained $\delta$-ISS-CLF satisfying the incremental stability conditions under the controller $g_{\bar{\theta},\bar{b}}$.
@@
Consider a dt-CS $\Xi$ given in (1). Let $V_{\theta,b}$ be the neural $\delta$ISS control Lyapunov function and let $g_{\bar{\theta},\bar{b}}$ be the corresponding controller. Let the optimal value of SCP (9), $\eta_{S}^{*}$, be obtained using data samples collected from the state-space and input space. Then under assumptions 3 and 4, if
@@
where $\mathcal{L}=\max\{\sqrt{2}\mathcal{L}_{L}+2\mathcal{L}_{1},\sqrt{2}\mathcal{L}_{L}+2\mathcal{L}_{2},\sqrt{2}\mathcal{L}_{L}(\mathcal{L}_{x}+\sqrt{2}\mathcal{L}_{u}\mathcal{L}_{C}+1)+2(\mathcal{L}_{3}+\mathcal{L}_{w}),\mathcal{L}_{h}(\mathcal{L}_{x}+\sqrt{2}\mathcal{L}_{u}\mathcal{L}_{C}+k_{h})\}$, then the obtained $\delta$-ISS control Lyapunov function upon solving the SCP ensures that the closed-loop system is incrementally input-to-state stable under the action of the controller $g_{\bar{\theta},\bar{b}}$.
Proof 3.3
The proof is presented in Appendix B.
3.3 Training Procedure of the Neural $\delta$-ISS-CLF and Corresponding Neural Controller
Here, we utilize the derived condition (10) from Section 3.2 and propose a training framework to synthesize provably correct $\delta$-ISS-CLF and the controller parametrized as neural networks. In particular, we train $\delta$-ISS-CLF and controller simultaneously to achieve formal guarantees on their validity by constructing suitable loss functions that incorporate the satisfaction of conditions (6a)-(6d) and (10) over the state-space and input-space.
We consider (9a)-(9d) as sub-loss functions to construct the actual loss function. The sub-loss functions are:
$L_{0}(\psi,\eta)$ $=\sum_{x,\hat{x}\in\mathcal{X}}\max\big{(}0,(-V_{\theta,b}(x,\hat{x})+k_{1}|x-\hat{x}|^{\gamma_{1}}-\eta)\big{)},$ (11a)
$L_{1}(\psi,\eta)$ $=\sum_{x,\hat{x}\in\mathcal{X}}\max\big{(}0,(V_{\theta,b}(x,\hat{x})-k_{2}|x-\hat{x}|^{\gamma_{2}}-\eta)\big{)},$ (11b)
$L_{2}(\psi,\eta)$ $=\sum_{x,\hat{x}\in\mathcal{X},w,\hat{w}\in{\mathcal{W}}}\max\big{(}0,(V_{\theta,b}(f(x,g_{\bar{\theta},\bar{b}}(x,w)),f(\hat{x},g_{\bar{\theta},\bar{b}}(\hat{x},\hat{w})))-V_{\theta,b}(x,\hat{x})+k_{3}|x-\hat{x}|^{\gamma_{3}}$
$\quad-k_{w}|w-\hat{w}|^{\gamma_{w}}-\eta)\big{)},$ (11c)
$L_{3}(\psi,\eta)$ $=\sum_{x\in\mathcal{X}}\max\big{(}0,h(f(x,g_{\bar{\theta},\bar{b}}(x,w)))-k_{h}h(x)-\eta\big{)},$ (11d)
where $\psi=[\theta,b,\bar{\theta},\bar{b}]$ and $\eta$ are trainable parameters. As mentioned, the actual loss function is a weighted sum of the sub-loss functions and is denoted by
$L(\psi,\eta)=c_{0}L_{0}(\psi,\eta)+c_{1}L_{1}(\psi,\eta)+c_{2}L_{2}(\psi,\eta)+c_{3}L_{3}(\psi,\eta),$ (12)
where $c_{0},c_{1},c_{2},c_{3}\in{\mathbb{R}}^{+}$ are the weights of the sub-loss functions $L_{0}(\psi,\eta),L_{1}(\psi,\eta),L_{2}(\psi,\eta)$, and $L_{3}(\psi,\eta)$, respectively.
Next, the Lipschitz continuity of the candidate neural networks following Assumption 4 needs to be satisfied. Since neural network candidates consist of slope-restricted activation layers, the assumption is already satisfied. Now, the following lemma addresses the guarantee of the Lipschitz bound over the training procedure.

### v2

@@ 2.3 Problem Formulation
Problem 2.10 can be reformulated as finding the $\delta$-ISS-CLF function that satisfies the conditions of Definition 2.6 under the existence of some forward invariant controller $g$.
@@ 3 Neural $\delta$-ISS Control Lyapunov Function
${\forall x,\hat{x}\in{\mathbf{X}},x=\hat{x}:V(x,\hat{x})=0,}$ (6a)
$\forall x,\hat{x}\in{\mathbf{X}},x\neq\hat{x}:-V(x,\hat{x})+\alpha_{1}(|x-\hat{x}|)\leq\eta,$ (6b)
$\forall x,\hat{x}\in{\mathbf{X}},x\neq\hat{x}:V(x,\hat{x})-\alpha_{2}(|x-\hat{x}|)\leq\eta,$ (6c)
$\forall x,\hat{x}\in{\mathbf{X}},x\neq\hat{x},\forall w,\hat{w}\in{\mathbf{W}}:$
$\quad V(f(x,g(x,w)),f(\hat{x},g(\hat{x},\hat{w})))-V(x,\hat{x})+\alpha_{3}(|x-\hat{x}|)-\sigma(|w-\hat{w})|)\leq\eta,$ (6d)
$\forall x\in{\mathbf{X}},\forall w\in{\mathbf{W}}:\ {h(f(x,g(x,w)))-h(x)}\leq\eta,$ (6e)
@@
Note that $V(\cdot,\cdot)\in\{V|V:{\mathbf{X}}\times{\mathbf{X}}\rightarrow{\mathbb{R}}_{0}^{+}\},g(\cdot,\cdot)\in\{g|g:{\mathbf{X}}\times{\mathbf{W}}\rightarrow{\mathbf{U}}\},\alpha_{1},\alpha_{2},\alpha_{3}\in\mathcal{K}_{\infty},\sigma\in\mathcal{K}$. If the optimal solution of ROP $\eta^{*}\leq 0$, under the constraints defined in (6) leads to the satisfaction of conditions of Definition 2.6 and 2.8, and thus the function V(x,x^)V($\hat{x}$) will be a valid $\delta$-ISS-CLF for the unknown system enforcing the system to be $\delta$-ISS under the forward invariant controller $g$.
@@
    The function $f$ is unknown, therefore we cannot directly incorporate conditions (6d) and (6e). So, solving the ROP becomes non-trivial.
@@
    The structures of the $\mathcal{K}_{\infty}$ functions $\alpha_{1},\alpha_{2},\alpha_{3}$ and class $\mathcal{K}$ function $\sigma$ are unknown.
@@
To overcome these challenges, the following subsections made some assumptions. To address challenge (C1), we parametrize $\delta$-ISS-CLF and the controller as feed-forward neural networks denoted by $V_{\theta,b}$ and $g_{\bar{\theta},\bar{b}}$, respectively, where $\theta,\bar{\theta}$ are weight matrices and $b,\bar{b}$ are bias vectors.
3.1 Construction of SOP and formal verification
The $\delta$-ISS-CLF neural network consists of an input layer with $2n$ (two times the dimension of the system $n$) and an output layer with one neuron, reflecting the scalar nature of $\delta$-ISS-CLF. The network includes $l_{v}$ hidden layers with each hidden layer containing $h_{v}^{i},i\in[1;l_{v}]$ neurons, where both values are arbitrarily chosen.
The activation function of all the layers except the output layer is chosen to be any slope-restricted nonlinear function $\varphi(\cdot)$ (for example, ReLU, Sigmoid, Tanh, etc.). Hence, the resulting neural network function is obtained by recursively applying the activation function as follows:
$\begin{cases}t^{0}=[x^{\top},\hat{x}^{\top}]^{\top},x,\hat{x}\in{\mathbf{X}},\\ t^{i+1}=\phi_{i}(\theta^{i}t^{i}+b^{i})\hskip 1.99997pt\text{for}\hskip 1.99997pti\in[0;l_{v}-1],\\ V_{\theta,b}(x,\hat{x})=\theta^{l_{v}}t^{l_{v}}+b^{l_{v}},\end{cases}$
@@ 3.1 Construction of SOP and formal verification
Now, to overcome the challenges (C2) and (C3), we raise the following assumptions:
@@
We consider having access to the black box or simulator model of the system. Hence, given a state-input pair $(x,u)$, we will be able to generate the next state $f(x,u)$.
@@
We assume that the class $\mathcal{K}_{\infty}$ functions $\alpha_{i},i\in\{1,2,3\}$ are of degree $\gamma_{i}$ with respect to |xx^||x-$\hat${x}| and the class $\mathcal{K}$ function $\sigma$ is of degree $\gamma_{w}$ with respect to |ww^||w-$\hat${w}|, i.e., $\alpha_{i}(|x-\hat{x}|)=k_{i}|x-\hat{x}|^{\gamma_{i}}$, $i\in\{1,2,3\}$ and $\sigma(|w-\hat{w}|)=k_{w}|w-\hat{w}|^{\gamma_{w}}$, where $k:=[k_{1},k_{2},k_{3},k_{w}]$ and $\Gamma:=[\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{w}]$ are user-defined parameters. The choice of $\Gamma$ is made in such a way that these functions become convex.
Now, for training the neural $\delta$-ISS-CLF and the controller, one requires data from the state space and the input space. Also, to overcome the challenge C4, we use a finite number of samples from the compact state space $\mathbf{X}$} and the input space $\mathbf{W}$}. To do so, we collect $N$ samples $x_{s}$ from the state space $\mathbf{X}$} where $s\in[1;N]$. We consider a ball $B_{\varepsilon_{x}}(x_{s})$ around each sample $x_{s}$ with radius $\varepsilon_{x}$ such that for all $x\in{\mathbf{X}}$, there exists an $x_{s}$ such that $|x-x_{s}|\leq\varepsilon_{x}$. This ensures that, $\bigcup_{s=1}^{N}B_{\varepsilon_{x}}(x_{s})\supset{\mathbf{X}}$. A similar strategy has been applied to collect $M$ samples $w_{p}$ from the input space $\mathbf{W}$}, where the radius of the ball is $\varepsilon_{u}$.
@@
$\mathcal{X}\!=\!\{x_{s}|\bigcup_{s=1}^{N}B_{\varepsilon_{x}}(x_{s})\!\supset\!{\mathbf{X}}\},\ {\mathcal{W}}\!=\!\{w_{p}|\bigcup_{p=1}^{M}B_{\varepsilon_{u}}(w_{p})\!\supset\!{\mathbf{W}}\}.$ (7)
@@
${\forall x_{q},x_{r}\in\mathcal{X},x_{q}=x_{r}:V_{\theta,b}(x_{q},x_{r})=0,}$ (8a)
$\forall x_{q},x_{r}\in\mathcal{X},x_{q}\neq x_{r}:-V_{\theta,b}(x_{q},x_{r})\!+\!k_{1}|x_{q}-x_{r}|^{\gamma_{1}}\!\leq\!\eta,$ (8b)
$\forall x_{q},x_{r}\in\mathcal{X},x_{q}\neq x_{r}:V_{\theta,b}(x_{q},x_{r})-k_{2}|x_{q}-x_{r}|^{\gamma_{2}}\leq\eta,$ (8c)
$\forall x_{q},x_{r}\in\mathcal{X},x_{q}\neq x_{r},\forall w_{q},w_{r}\in{\mathcal{W}}:$
$\quad V_{\theta,b}(f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q})),f(x_{r},g_{\bar{\theta},\bar{b}}(x_{r},w_{r})))-V_{\theta,b}(x_{q},x_{r})+k_{3}|x_{q}-x_{r}|^{\gamma_{3}}-k_{w}|w_{q}-w_{r}|^{\gamma_{w}}\leq\eta,$ (8d)
$\forall x_{q}\!\in\!\mathcal{X},w_{q}\in{\mathcal{W}}:{h(f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q})))\!-\!h(x_{q})}\!\leq\!\eta.$ (8e)
Note that the optimization program becomes convex as all the constraints, as well as the objective function, are convex with respect to the decision variable. Now, given the finite number of data samples, the SCP involves a finite set of constraints, making its solution computationally tractable. Let $\eta_{S}^{*}$ denote the optimal solution of the SCP. To demonstrate that the solution of the SCP is also a feasible solution for the proposed ROP, we impose the following assumptions regarding Lipschitz continuity:
@@
We assume the Lipschitz constants of the dynamics are known, even if the system is unknown. Additionally, these constants can also be estimated following a similar procedure in [25, Algorithm 2].
@@
We assume that the candidate neural $\delta$-ISS-CLF is Lipschitz continuous with Lipschitz bound $\mathcal{L}_{L}$ with respect to (x,x^)($\hat{x}$) over the set $\mathbf{X}$}. Similarly, the controller neural network has a Lipschitz bound $\mathcal{L}_{C}$.
@@
The class $\mathcal{K}_{\infty}$ functions and the class $\mathcal{K}$ function of Definition 2.6 are Lipschitz continuous with Lipschitz constants $\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}$, and $\mathcal{L}_{w}$, respectively by virtue of Assumption 2, as it is evident that convex continuous functions are Lipschitz continuous inside a compact set. One can estimate the values using the values of $k$ and $\Gamma$. In addition, the Lipschitz constant $\mathcal{L}_{h}$ of the function $h$ is already predefined due to the known structure of $h$.
Under Assumptions 3 and 4, the following theorem outlines the connection of the solution of SCP (8) to that of ROP (6), providing a formal deterministic guarantee to the obtained $\delta$-ISS-CLF satisfying the incremental stability conditions under the controller $g_{\bar{\theta},\bar{b}}$.
@@
[…]
@@
where $\mathcal{L}=\max\{\sqrt{2}\mathcal{L}_{L}+2\mathcal{L}_{1},\sqrt{2}\mathcal{L}_{L}+2\mathcal{L}_{2},\sqrt{2}\mathcal{L}_{L}(\mathcal{L}_{x}+\sqrt{2}\mathcal{L}_{u}\mathcal{L}_{C}+1)+2(\mathcal{L}_{3}+\mathcal{L}_{w}),\mathcal{L}_{h}(\mathcal{L}_{x}+\sqrt{2}\mathcal{L}_{u}\mathcal{L}_{C}+1)\}$, then the obtained $\delta$-ISS control Lyapunov function upon solving the SCP ensures that the closed-loop system is incrementally input-to-state stable under the action of the controller $g_{\bar{\theta},\bar{b}}$ with a deterministic correctness guarantee.
Proof:
The proof is presented in Appendix B. $\square$
Remark 3.3
In this work, although we use a scenario-based sampling approach to approximate the ROP, our samples are generated deterministically, unlike the random sampling in [32] that yields probabilistic guarantees. This deterministic sampling, together with the Lipschitz-based validity condition, ensures the deterministic correctness of the synthesized controller and $\delta$-ISS-CLF.
3.2 Formulation of Loss functions and Training procedure
Here, we utilize the derived condition (9) from Section 3.1 and propose a training framework to synthesize provably correct $\delta$-ISS-CLF and the controller parametrized as neural networks. In particular, we train $\delta$-ISS-CLF and controller simultaneously to achieve formal guarantees on their validity by constructing suitable loss functions that incorporate the satisfaction of conditions (6b)-(6e) and (9) over the state-space and input-space.
We consider (8a)-(8e) as sub-loss functions to construct the actual loss function. The sub-loss functions are:
$L_{0}(\psi,\eta)$ $),\displaystyle=\sum_{x=\hat{x}}\max\big(0,V_{\theta,b}(x,\hat{x}$$\big)$, (10a)
$L_{1}(\psi,\eta)$ $=\sum_{x\neq\hat{x}}\max\big(0,(-V_{\theta,b}(x,\hat{x})+k_{1}|x-\hat{x}|^{\gamma_{1}}-\eta)\big),$ (10b)
$L_{2}(\psi,\eta)$ $=\sum_{x\neq\hat{x}}\max\big(0,(V_{\theta,b}(x,\hat{x})-k_{2}|x-\hat{x}|^{\gamma_{2}}-\eta)\big),$ (10c)
$L_{3}(\psi,\eta)$ $=\sum_{x\neq\hat{x}}\max\big(0,(V_{\theta,b}(f(x,g_{\bar{\theta},\bar{b}}(x,w)),f(\hat{x},g_{\bar{\theta},\bar{b}}(\hat{x},\hat{w})))-V_{\theta,b}(x,\hat{x})+k_{3}|x-\hat{x}|^{\gamma_{3}}-k_{w}|w-\hat{w}|^{\gamma_{w}}-\eta)\big),$ (10d)
$L_{4}(\psi,\eta)$ $=\sum\max\big(0,h(f(x,g_{\bar{\theta},\bar{b}}(x,w)))-k_{h}h(x)-\eta\big),$ (10e)
where, $x,\hat{x}\in\mathcal{X},w,\hat{w}\in\mathcal{W}$, with $\psi=[\theta,b,\bar{\theta},\bar{b}]$ and $\eta$ being the trainable parameters. As mentioned, the actual loss function is a weighted sum of the sub-loss functions and is denoted by
$L(\psi,\eta)$ $=c_{0}L_{0}(\psi,\eta)+c_{1}L_{1}(\psi,\eta)+c_{2}L_{2}(\psi,\eta)+c_{3}L_{3}(\psi,\eta)+c_{4}L_{4}(\psi,\eta),$ (11)
where $c_{0},c_{1},c_{2},c_{3},c_{4}\in{\mathbb{R}}^{+}$ are the weights of the sub-loss functions $L_{0}(\psi,\eta),L_{1}(\psi,\eta),L_{2}(\psi,\eta),L_{3}(\psi,\eta)$ and $L_{4}(\psi,\eta)$ respectively.
Next, the Lipschitz continuity of the candidate neural networks following Assumption 4 needs to be satisfied. Since neural network candidates consist of Lipschitz activation layers, the assumption is already satisfied. Now, the following lemma addresses the guarantee of the Lipschitz bound over the training procedure.